# Multiview Random Vector Functional Link Network for Predicting DNA-Binding Proteins

### v1

@@ 5 Proposed Multiview Random Vector Functional Link (MvRVFL) Network
This section offers an in-depth explanation of the proposed MvRVFL model. Initially, we outline the generic mathematical framework of the proposed MvRVFL model, specifically tailored to handle data originating from two distinct views. During training on one view, the influence of other views is incorporated by introducing a coupling term into the primal form of the proposed optimization problem. An intuitive illustration of the MvRVFL model is shown in Fig. 4. Let $Z_{1}$ and $Z_{2}$ represent the nonlinear projection of the class samples corresponding to Vw-$A$ and Vw-$B$, as defined by: $Z_{1}=[X_{1}\hskip 5.69046ptH_{1}]$ and $Z_{2}=[X_{2}\hskip 5.69046ptH_{2}]$. Here $X_{1}$ and $X_{2}$ represent the input matrix corresponding to Vw-$A$ and Vw-$B$, and $H_{1}$ and $H_{2}$ denote the hidden layer matrix corresponding to Vw-$A$ and Vw-$B$, respectively. These matrices are obtained by transforming $X_{1}$ and $X_{2}$ using randomly initialized weights and biases and then applying a nonlinear activation function. The target matrix is denoted by $Y$. The proposed optimization problem of MvRVFL model is articulated as follows:
@@
    The terms $\|\beta_{1}\|$ and $\|\beta_{2}\|$ are regularization components for Vw-$A$ and Vw-$B$, respectively. These terms are employed to mitigate overfitting by constraining the capacities of the classifier sets for both views.
@@
    The primal optimization function comprises two distinct classification objectives for each view, linked by the coupling term $\rho\xi_{1}^{t}\xi_{2}$. Here, $\rho$ represents an additional regularization constant, referred to as the coupling parameter. This term serves to minimize the product of the error variables for both views as well as trade-off parameters between both views.
@@
Using the Karush-Kuhn-Tucker (K.K.T.) conditions, we have
@@
After computing the optimal values of $\beta_{1}$ and $\beta_{2}$, the classification of a new input data point $x$ into either the $+1$ or $-1$ class can be determined as follows:
@@
    $class^{A}(x^{A})=\underset{i\in\{1,2\}}{\arg\max}\{y_{A_{i}}\},$ (30)
    $\text{where}\hskip 5.69046pty_{A}=[x^{A}\hskip 5.69046pt\phi(x^{A}W^{A}+b^{A})]\beta_{1}\hskip 2.84544pt\text{and}\hskip 2.84544pty_{A}=(y_{A_{1}},y_{A_{2}})$
    and
    $class^{B}(x^{B})=\underset{i\in\{1,2\}}{\arg\max}\{y_{B_{i}}\},$ (31)
    $\text{where}\hskip 5.69046pty_{B}=[x^{B}\hskip 5.69046pt\phi(x^{B}W^{B}+b^{B})]\beta_{2}\hskip 2.84544pt\text{and}\hskip 2.84544pty_{B}=(y_{B_{1}},y_{B_{2}}).$
@@
    $class(x)=\underset{i\in\{1,2\}}{\arg\max}\{y_{c_{i}}\},$ (32)
    where
    $y_{c}=\frac{1}{2}\left([x^{A}\hskip 5.69046pt\phi(x^{A}W^{A}+b^{A})]\beta_{1}+[x^{B}\hskip 5.69046pt\phi(x^{B}W^{B}+b^{B})]\beta_{2}\right)\hskip 2.84544pt\text{and}\hskip 2.84544pty_{c}=(y_{c_{1}},y_{c_{2}}).$
Here, $W^{A}$ ($W^{B}$) and $b^{A}$ ($b^{B}$) are the randomly generated weights and biases corresponding to Vw-$A$ (Vw-$B$), respectively.
Remark: The proposed MvRVFL model is referred to as MvRVFL-1 if the classification of the test sample is determined by the functions outlined in Eq. (32). Also, we employ majority voting to determine the final anticipated output of the proposed MvRVFL model by aggregating predictions from Eqs. (30), (31) and (32), and we refer to it as MvRVFL-2 model. In the majority voting technique, the class label with the highest number of votes is selected as the final prediction.
The algorithm of the proposed MvTPSVM model is briefly described in Algorithm 1. Furthermore, evaluating the generalization performance of MvRVFL requires a theoretical examination of its generalization error. To achieve this, we apply Rademacher’s complexity theory. In the Appendix (Section A.1), we delve into the consistency error bound and the generalization error bound of the MvRVFL model.
@@ 5.1 Complexity Analysis of the Proposed MvRVFL Model
Let $(X_{1},X_{2},Y)$ denote the training set, where $X_{1}\in\mathbb{R}^{n\times m_{1}}$, $X_{2}\in\mathbb{R}^{n\times m_{2}}$ and $Y\in\mathbb{R}^{n\times 2}$. Here, $m_{1}$ and $m_{2}$ denote the number of features corresponding to Vw-$A$ and Vw-$B$ and $n$ represents the number of samples, respectively. In RVFL-based models, the complexity is governed by the necessity to compute matrix inverses for optimizing the output layer weights. Thus, the size of the matrices requiring inversion dictates the model’s complexity. Therefore, the time complexity of the RVFL model is $\mathcal{O}(n^{3})$ or $\mathcal{O}((m+h_{l})^{3})$, where $h_{l}$ represents the number of hidden nodes. Hence, the proposed MvRVFL model is required to inverse the matrix of dimension $(m_{1}+m_{2}+2h_{l})\times(m_{1}+m_{2}+2h_{l})$. Therefore, the time complexity of the MvRVFL model is $\mathcal{O}((m_{1}+m_{2}+2h_{l})^{3})$.
@@
Require: $X_{1}\in\mathbb{R}^{n\times m_{1}}$ and $X_{2}\in\mathbb{R}^{n\times m_{2}}$ be the input matrix of Vw-$A$ and Vw-$B$, respectively and $Y\in\mathbb{R}^{n\times 2}$ be the target matrix. Here, $n$ represents the number of samples, while $m_{1}$ and $m_{2}$ represent the number of features of the input sample corresponding to Vw-$A$ and Vw-$B$, respectively.
1:$W^{A}\in\mathbb{R}^{m_{1}\times h_{l}}$, $b^{A}\in\mathbb{R}^{n\times h_{l}}$, $W^{B}\in\mathbb{R}^{m_{2}\times h_{l}}$, and $b^{B}\in\mathbb{R}^{n\times h_{l}}$ are randomly initialized corresponding to Vw-$A$ and Vw-$B$, with all columns of $b^{A}$ and $b^{B}$ being identical.

### v2

@@ 4 Proposed Multiview Random Vector Functional Link (MvRVFL) Network
This section offers an in-depth explanation of the proposed MvRVFL model. Initially, we outline the generic mathematical framework of the proposed MvRVFL model, specifically tailored to handle data originating from two distinct views. While training on a single view, the impact of additional views is taken into account by adding a coupling component to the primal formulation of the optimization problem being proposed. An intuitive illustration of the MvRVFL model is shown in Fig. 3. Let $Z_{1}$ and $Z_{2}$ represent the nonlinear projection of the class samples corresponding to Vw-$A$ and Vw-$B$, as defined by: $Z_{1}=[X_{1}\hskip 5.69046ptH_{1}]$ and $Z_{2}=[X_{2}\hskip 5.69046ptH_{2}]$. Here, $H_{1}$ and $H_{2}$ denote the hidden layer matrix corresponding to Vw-$A$ and Vw-$B$, respectively. These matrices are obtained by transforming $X_{1}$ and $X_{2}$ using randomly initialized weights and biases and then applying a nonlinear activation function. The optimization formulation for the proposed MvRVFL model is presented as follows:
@@
    The terms $\|\beta_{1}\|$ and $\|\beta_{2}\|$ are regularization components for Vw-$A$ and Vw-$B$, respectively. These elements help prevent overfitting by limiting the complexity of the classifier sets associated with each specific view.
@@
    The primal optimization function consists of two separate classification goals for each view, connected through the coupling term $\rho\xi_{1}^{t}\xi_{2}$. In this context, $\rho$ is an extra regularization constant, known as the coupling parameter. This term aims to reduce the product of the error variables from both views while balancing the trade-off parameters between them.
@@
By utilizing the Karush-Kuhn-Tucker (KKT) conditions, we derive the following:
@@
Once the optimal values of $\beta_{1}$ and $\beta_{2}$ have been determined, a new data point $x$ can be classified by using the following equation:
@@
    $class^{A}(x^{A})=sign(y_{A}),\penalty 10000\ \text{where}\hskip 5.69046pty_{A}=[x^{A}\hskip 5.69046pt\phi(x^{A}W^{A}+b^{A})]\beta_{1},\penalty 10000\ \text{and}$ (23)
    $class^{B}(x^{B})=sign(y_{B}),\penalty 10000\ \text{where}\hskip 5.69046pty_{B}=[x^{B}\hskip 5.69046pt\phi(x^{B}W^{B}+b^{B})]\beta_{2}.$ (24)
@@
    $class(x)=sign(y_{c}),\penalty 10000\ \text{where}\penalty 10000\ \penalty 10000\ y_{c}=\frac{1}{2}\left(y_{A}+y_{B}\right).$ (25)
Here, $W^{A}$ ($W^{B}$) and $b^{A}$ ($b^{B}$) are randomly generated weights and biases for Vw-$A$ (Vw-$B$), respectively.
Remark: The proposed MvRVFL model is referred to as MvRVFL-1 if the classification of the test sample is determined by the functions outlined in Eq. (25). Also, we employ majority voting to determine the final anticipated output of the proposed MvRVFL model by aggregating predictions from Eqs. (23), (24) and (25), and we refer to it as MvRVFL-2 model. In the majority voting approach, the class label with the highest number of votes is selected as the final prediction.
@@ 4.1 Complexity Analysis of the Proposed MvRVFL Model
Let $(X_{1},X_{2},Y)$ represent the training dataset, where $X_{1}\in\mathbb{R}^{n\times m_{1}}$, $X_{2}\in\mathbb{R}^{n\times m_{2}}$, and $Y\in\mathbb{R}^{n\times 1}$. Here, $m_{1}$ and $m_{2}$ refer to the number of features for Vw-$A$ and Vw-$B$, respectively, while $n$ indicates the total number of samples. In RVFL-based models, the computational complexity arises from the need to compute matrix inverses to optimize the weights of the output layer. Thus, the size of the matrices requiring inversion dictates the model’s complexity. Therefore, the time complexity of the RVFL model is $\mathcal{O}(n^{3})$ or $\mathcal{O}((m+h_{l})^{3})$, where $h_{l}$ represents the number of hidden nodes. Hence, the proposed MvRVFL model is required to inverse the matrix of dimension $(m_{1}+m_{2}+2h_{l})\times(m_{1}+m_{2}+2h_{l})$. As a result, the time complexity of the MvRVFL model is $\mathcal{O}((m_{1}+m_{2}+2h_{l})^{3})$. The procedure for the MvTPSVM model is outlined concisely in Algorithm 1.
@@
Require: $X_{1}\in\mathbb{R}^{n\times m_{1}}$ and $X_{2}\in\mathbb{R}^{n\times m_{2}}$ be the input matrix of Vw-$A$ and Vw-$B$, respectively and $Y\in\mathbb{R}^{n\times 1}$ be the target vector. Here, $n$ represents the number of samples, while $m_{1}$ and $m_{2}$ represent the number of features of the input sample corresponding to Vw-$A$ and Vw-$B$, respectively.
[…]